JET-COOLED SPECTROSCOPY OF THE A^{6}Σ^{+} − X^{6}Σ^{+} SYSTEM OF CHROMIUM MONOHYDRIDE

PRADYOT K. CHOWDHURY, ANTHONY J. MERER and \underline{SCOTT J. RIXON}, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver B.C. V6T 1Z1, Canada.

Astrophysical interest in CrH has motivated several studies of its A^{6}Σ^{+} − X^{6}Σ^{+} system^{a}. In this work, pulsed laserinduced fluorescence spectra of the (v′, 0) bands with v′ = 1, 2 and 3 have been recorded under supersonic jet-cooled conditions following reaction of laser-ablated chromium atoms with methanol. Because of the large rotational constant and the jet-cooling, only lines with N′′ ≤ 5 are appreciably intense. Satellite branches with ΔN ≠ ΔJ, normally forbidden by Hund’s case (b) selection rules, readily appear in all three bands as previously observed for the (0,0) band^{a}. Lines of the (2,0) and (3,0) bands are assigned here for the first time; strong perturbations considerably complicate the rotational structures of these bands.

aS. Shin, D. J. Brugh and M. D. Morse, Astrophys. J. 619, 407−411 (2004) and references therein.